Alonzo bought a sofa on sale for $123.20 . This price was 20% of the original price.
What was the original price?

Answer by pedjajov(51)

Since price for sofa (s) is 20% of the original price (p) it is:
s=20% of p
s=p*(20/100)
s=p/5
p=5s = 5*123.20 = 616 dollars was original price